OpenStudy (anonymous):

Chris and Justine are purchasing a townhouse and finance $112,900 with a 25-year 6/3 ARM at 4.85% with a 2/12 cap structure. What will their payments be at the beginning of the seventh year assuming they are charged the maximum interest rate for that year? $759.70 $557.09 $674.48 $650.17

OpenStudy (anonymous):

@amistre64 I used this formula ->Bm = P[((1+i)^(n*t)-(1-i)^(m))/((1+i)^(n*t)-(1))] I plugged in P=112900 I=4.85%/12=0.004041667 N=12 T=25 M=7*12=84 I got http://www.wolframalpha.com/input/?i=112900%5B%28%281%2B0.004041667%29%5E%2812*25%29-%281-0.004041667%29%5E%2884%29%29%2F%28%281%2B0.004041667%29%5E%2812*25%29-%281%29%29%5D Then I add new info to it I=0.0685/12 = 0.005708333 <--- ( 4.85 + 2 = 6.85%) N=12 T-25-6 = 19 and I got http://www.wolframalpha.com/input/?i=112900%5B%28%281%2B0.005708333%29%5E%2812*19%29-%281-0.005708333%29%5E%2884%29%29%2F%28%281%2B0.005708333%29%5E%2812*19%29-%281%29%29%5D

OpenStudy (anonymous):

I dont know what to do next..

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@ganeshie8